# Learning Personalized Decision Support Policies

### v2

@@ Abstract
Individual human decision-makers may benefit from different forms of support to improve decision outcomes, but when each form of support will yield better outcomes? In this work, we posit that personalizing access to decision support tools can be an effective mechanism for instantiating the appropriate use of AI assistance. Specifically, we propose the general problem of learning a decision support policy that, for a given input, chooses which form of support to provide to decision-makers for whom we initially have no prior information. We develop Modiste, an interactive tool to learn personalized decision support policies. Modiste leverages stochastic contextual bandit techniques to personalize a decision support policy for each decision-maker and supports extensions to the multi-objective setting to account for auxiliary objectives like the cost of support. We find that personalized policies outperform offline policies, and, in the cost-aware setting, reduce the incurred cost with minimal degradation to performance. Our experiments include various realistic forms of support (e.g., expert consensus and predictions from a large language model) on vision and language tasks. Our human subject experiments validate our computational experiments, demonstrating that personalization can yield benefits in practice for real users, who interact with Modiste.
Introduction
@@
To improve decision outcomes, human decision-makers can use various forms of support to inform their opinions before making a final decision (Keen 1980). Decision-makers with differing expertise may benefit from different forms of support on a given input. For example, one radiologist may provide a better diagnosis of a chest X-ray by leveraging model predictions (Kahn Jr 1994) while another may perform better after viewing suggestions from senior radiologists (Briggs et al. 2008): see Figure 1. In this paper, we study how to improve decision outcomes by personalizing which form of support we provide to a decision-maker on a case-by-case basis.
@@
Figure 1: Depending on the input, decision-makers need different forms of decision support to make correct decisions. Modiste personalizes access to the right form of support at the right time for the right decision-maker. Here, Alice would not benefit from model access, while Bob would not benefit from a senior consult.
@@
Since artificial intelligence (AI) is increasingly used as a form of decision support (Lai et al. 2023), responsible machine learning (ML) model deployment requires clarity on who should have access to model outputs and when model outputs can be safely exposed to decision-makers (Amodei et al. 2016). Regulation increasingly calls for the “effective and appropriate use” of AI (Biden 2023), requiring careful consideration of when models ought to be accessible to decision-makers. In this paper, we formalize learning a decision support policy that dictates for each individual decision-maker when additional support (e.g., LLM output) should be viewed and used for a given input.
@@
While prior work has assumed access to offline human decisions under support (Laidlaw and Russell 2021; Charusaie et al. 2022) or oracle queries of human behavior (De-Arteaga, Dubrawski, and Chouldechova 2018; Mozannar and Sontag 2020) to learn decision support policies, we argue that this data is unrealistic to obtain in practice across all available forms of support for a new decision-maker. Thus, for individuals for whom we have no prior information initially, we propose learning how to personalize support online. We develop Modiste, [^1] While a “modiste” usually refers to someone who tailors clothing and makes dresses/hats, we use the term to capture our tool’s ability to alter a policy to a decision-maker. an interactive tool that leverages off-the-shelf stochastic contextual bandit algorithms  (Li et al. 2010) to learn decision support policies. Since providing support may be costly (Arbiser, Folpe, and Weiss 2001; Mimra, Rasch, and Waibel 2016), we incorporate a second objective, the cost of support, and implement mechanisms for trading-off performance and cost when using Modiste.
@@
Our computational experiments explore the utility of personalization in both the cost-agnostic and cost-aware settings. Based on these experiments, we characterize decision-maker expertise into profiles where personalized policies outperform offline ones. We demonstrate that if there is no benefit of personalization, Modiste recovers fixed policies, which always provide one form of support. In the cost-aware setting, we show that personalization can reduce the incurred cost with minimal degradation to performance.
@@
To validate Modiste on real users ($N=125$), we conduct human subject experiments, where we explore forms of support that include expert consensus, outputs from an LLM, or predictions from a classification model. In contrast to prior work that only tests offline policies or evaluates in simulation, we demonstrate how Modiste can be used to learn personalized decision support policies online on both vision and language tasks. We emphasize our main contributions:
@@
1. Formalizing decision support policies. We propose a formulation for learning a personalized decision support policy that selects the form of support that maximizes a given decision-maker’s performance. We introduce Modiste, a tool to instantiate our formulation using existing methods from stochastic contextual bandits, and provide a multi-objective extension to incorporate the cost of support. We open-source Modiste as a tool to encourage the adoption of personalized decision support policies.
@@
2. Evaluating personalized policies in realistic settings. We demonstrate the importance of online learning to personalize policies for new decision-makers through both computational and human subject experiments on vision and language tasks. We characterize under which settings we would expect personalized policies to improve performance and/or cost. Our human subject experiments, where real users interact with Modiste, validate our findings from the computational experiments on synthetic decision-makers, demonstrating the benefits of personalized policies in practice.
Related Work
@@
Our study of decision support policies has implications for safely deploying ML models to interact with users. This topic is of heightened importance, particularly in light of recent calls for the “effective and appropriate use” of AI in US President Biden’s Executive Order (Biden 2023) and for clarity on when to “decide not to use [an] AI system” per the EU AI Act (EUA 2023). The disuse of AI can caution downstream misuse of models for assistive decision-making (Brundage et al. 2018). The refusal to use AI assistance can be strategic to empower decision-makers, thus preventing their overreliance on models and encouraging their agency on the task at hand (Gordon and Mugar 2020; Barabas 2022). Each decision-maker may require a different level of use to promote effective use of AI assistance in their decision-making (Kirk et al. 2024); for instance, experts and novices may prefer LLM access in different settings on theorem proving tasks (Collins et al. 2023b). Our experiments engage with such settings by learning when to provide LLM support for language-based tasks via a personalized policy.
@@
While various forms of decision support have been proposed, such as expert consensus (Scheife et al. 2015) and changes to machine interfaces (Roda 2011), more recent forms of support focus on algorithmic tools where decision-makers are aided by machine learning (ML) models (Phillips-Wren 2012; Gao et al. 2021; Bastani, Bastani, and Sinchaisri 2022). In some prior work, the human does not always make the final decision, such as those that learn to defer decisions from a model to a single decision-maker (Madras, Pitassi, and Zemel 2018; Mozannar and Sontag 2020) or others that jointly learn an allocation function between a model and a pool of decision makers (Keswani, Lease, and Kenthapadi 2021; Hemmer et al. 2022). In our setting, the human is always the decision-maker, even after viewing any form of support: this reflects many real-world applications. This setting includes ones where humans make the final decisions with support from ML models (Green and Chen 2019; Lai et al. 2023), as well as ones where humans make decisions when provided with additional information beyond a model prediction, e.g., explanations (Bansal et al. 2021b), uncertainty (Zhang, Liao, and Bellamy 2020), conformal sets (Babbar, Bhatt, and Weller 2022). While these studies always show a single form of support, recent works consider adapting when AI support is shown to users with a fixed policy. Ma et al. (2023) fit a decision tree to offline user’s decisions to decide when to show AI support to users, and Buçinca et al. (2024) use offline reinforcement learning to estimate if AI support would be helpful, especially under time constraints (Swaroop et al. 2024). Our work considers general forms of support, beyond when to show AI support, and formalizes learning in which contexts each form of support should be provided to an unseen decision-maker online. An extensive comparison to prior work is in the Appendix.
@@
Figure 2: We illustrate the process of learning a decision support policy $\pi_{t}$ online to improve a decision-maker $h$’s performance. Without assuming access to prior offline data, our formulation learns a personalized policy online; each decision-maker’s learned policy may differ from that of another decision-maker if they have decisions ($\tilde{y}$) and thus different expertise. We also consider a cost-aware version of this learning problem that modifies how $a_{t}$ is selected.
Preliminaries
@@
A static decision support policy would not change throughout the interaction with the decision-maker. This reflects many prior user studies, per Lai et al. (2023); for example, decision-makers may always be presented with model predictions. We iteratively update $\pi$ at each time step to personalize the policy based on the decision-maker’s most recent prediction.
@@
Modiste: Learning Personalized Decision Support Policies
@@
The human decision-making process with various forms of support can be modeled as a stochastic contextual bandit problem, where the forms of support are the arms and ${\mathcal{X}}$ is the context space. In Algorithm 1, we outline an online algorithm for learning the policy $\pi^{*}$ that minimizes expected loss. In the cost-agnostic setting, our goal is to find an optimal decision support policy $\pi^{*}$ that minimizes $L_{h}(\pi)$. We can rewrite Eq. 1 as follows:
@@
where $r_{\textsc{A}_{i}}(x;h)=\mathbb{E}_{y|x}[\ell(y,h(x,\textsc{A}_{i}))]$ is the human prediction error for input $x$ and support $\textsc{A}_{i}$. Then, it can be shown that the optimal policy takes the form $\pi^{*}(x)=\operatorname*{arg\,min}_{\textsc{A}_{i}\in\mathcal{A}}r_{\textsc{A}_{i}}(x;h)$: see derivation in Appendix. For Modiste to run, we must maintain and update our estimate of human prediction error $r_{\textsc{A}_{i}}(x;h)$ (Step 1 in Algorithm 1) and of our policy $\pi$ (Step 2 in Algorithm 1)
@@
where $b_{\textsc{A}_{i},t}(x;h)$ corresponds to some exploration bonus. In the Appendix, we provide implementations of Modiste with LinUCB (Algorithm 2) and KNN (Algorithm 3).
@@
Extension to Cost-Aware Decision Support.
Providing decision support can be expensive (Arbiser, Folpe, and Weiss 2001; Mimra, Rasch, and Waibel 2016), so in addition to minimizing loss, it may be desirable to minimize the cost of providing support. We quantify the expected cost of a policy $\pi$ using $c(\pi)=\mathbb{E}_{x}\big{[}\sum_{i=1}^{k}{\pi(x)_{\textsc{A}_{i}}\cdot c(\textsc{A}_{i})}\big{]}$, where $c(\textsc{A}_{i})$ is the cost of providing support $\textsc{A}_{i}$. We reformulate our objective as $\pi^{*}_{\lambda}=\operatorname*{arg\,min}_{\pi\in\Pi}\lambda\cdot L_{h}(\pi)+(1-\lambda)\cdot c(\pi)$, where $\lambda\in[0,1]$. Since $c(\textsc{A}_{i})$ does not depend on $x$, it does not affect Step 1. Specifically, we can update Step 2 of Algorithm 1 to the following:
$\pi_{t+1}(x)=\operatorname*{arg\,min}_{\textsc{A}_{i}\in\mathcal{A}}\lambda\cdot\widehat{r}_{\textsc{A}_{i},t}(x;h)+(1-\lambda)\cdot c(\textsc{A}_{i})+b_{\textsc{A}_{i},t}(x;h).$
In practice, it may be unclear how to select $\lambda$, but one way to consider this problem may be to learn a decision support policy $\pi$ that achieves a certain level of performance at a minimal cost: given a tolerance threshold $\epsilon\in[0,1]$ for the expected loss, we can find a policy $\pi$ that achieves the minimum expected cost while maintaining the expected loss $L_{h}(\pi)$ within $\epsilon$ of the optimal loss $L_{h}^{\text{opt}}$, where $L_{h}^{\text{opt}}=\min_{\pi}L_{h}(\pi)$. Given a large number of interactions with a decision-maker $h$, we can simply run Algorithm 1 for multiple values of $\lambda$, i.e., sampling from $\lambda\in[0,1]$. From the set of corresponding policies learned for each $\lambda$, we can evaluate which $\lambda$ yields a policy with the lowest cost within $\epsilon$ of $L_{h}^{\text{opt}}$.
For a new decision-maker, it is often infeasible to extensively test many choices of $\lambda$ values simultaneously without repeating queries of the same input (e.g., as in our human subject experiments); thus, we leverage a hyper-parameter tuning strategy using population data. We follow the above workflow to find suitable values for $\lambda$ for each decision-maker in a population. From this set of candidate $\lambda$ values, we select a $\lambda$ to deploy by choosing the most common value (mode) of all $\lambda$ values that yield policies which get decision-makers within $\epsilon$ of their optimal loss. Although the decision-makers in the population may differ from the unseen decision-maker, we prefer this to a random selection of $\lambda$ at deployment.
@@
We provide an extendable interface for the study and deployment of decision-support policies. At each time step, Modiste sends each user’s predictions to a server running Algorithm 1, which identifies the next form of support for the next input. Modiste then updates the interface accordingly to reflect the selected form of support. Our tool can flexibly be linked with crowdsourcing platforms like Prolific (Palan and Schitter 2018). We implement three common forms of support: (1) Human Alone, where the human makes the decision solely based on the input, (2) Model Prediction, which shows decision-makers a model’s prediction for the given input (Bastani, Bastani, and Sinchaisri 2022), and (3) Expert Consensus, which presents the user with a distribution over labels from multiple annotators (Scheife et al. 2015). In Figure 3, we provide screenshots of the interface under each form of support. Participants are informed of their own correctness after each trial, as well as the correctness of the form of support (e.g., model prediction), if support was provided, so that participants can learn whether support ought to be relied upon.
Experimental Set-up
Before we validate Modiste, we overview the set-up of subsequent experiments, both computational and human subject. All other details are in the Appendix.
@@
In this task, we consider three forms of support: Human Alone, Model, Consensus. Our goal is to construct a setup reflecting a realistic setting in which different forms of support result in different strengths and weaknesses for decision-makers. To instantiate this setting, we deliberately corrupt images of different classes to evoke performance differences – necessitating that a decision-maker appropriately calibrate when to rely on each form of support. We consider $5$ of the animal classes in CIFAR-10; of these, we never corrupt images of Birds, do not corrupt images of Deers and Cats for the Model, and do not corrupt images of Horses and Frogs for the Consensus. We assume the cost of Human Alone is less than the cost of either form of support, and the cost of Model and Consensus is equal.
@@
The two forms of support are Human Alone and LLM, where the human is provided responses generated from InstructGPT3.5, text-davinci-003(Ouyang et al. 2022) using the same few-shot prompting scheme for MMLU as Hendrycks et al. (2020). We conducted pilot studies to select a subset of topics where the accuracy of the LLM and average human accuracy vary. We choose the following topics: Computer Science, US Foreign Policy, High School Biology, and Elementary Mathematics. The goal of this task is to evaluate whether we can learn personalized support “in-the-wild,” where we naturally expect people to excel at different topics, akin to real-world settings where decision-makers may have varying expertise. Again, the cost of Model is greater than the cost of Human Alone.
@@
  • Population-level, where the decision-maker receives a form of support based on the majority vote from 10 learned policies (breaking ties at random). For this baseline, the form of support may vary across contexts but is not personalized to individual needs. This baseline is akin to recent offline policy learning (Ma et al. 2023; Buçinca et al. 2024).
@@
Table 1: In the cost-agnostic setting, we evaluate the average excess loss $L_{h}(\pi)-L_{h}^{opt}$ (lower is better), and standard deviation across individuals in each expertise profile for both CIFAR-$3A$ (Left) and MMLU-$2A$ (Right). $L_{h}(\pi)$ is computed by averaging across the last 10 steps of 100 total time steps. We bold the variant with the lowest excess loss.
@@
Modiste-KNN $0.01\pm 0.01$ $\mathbf{0.05\pm 0.03}$ $\mathbf{0.05\pm 0.03}$
Table 2: In the cost-aware setting for MMLU-$2A$, we compare the expected loss $L_{h}(\pi)$ and the expected cost $c(\pi)$ (lower is better) averaged across the last 10 steps of 100 total time steps of 5 separate runs. We select two individuals who have a “varying” profile and two individuals who have a “strictly better” profile. We follow the described workflow to find a suitable $\lambda$, reported for each person in Table 5. We bold the algorithm that achieves the lowest cost within $\epsilon=0.05$ risk of each individual $h$’s $L_{h}^{\text{opt}}$.
Varying: $L_{h}^{\text{opt}}=0.13$
$r_{\textsc{H-Only}}=[0.3,0.1,0.5,0.3]$
$r_{\textsc{H-LLM}}=[0.5,0.1,0.0,0.1]$
Varying: $L_{h}^{\text{opt}}=0.10$
$r_{\textsc{H-Only}}=[0.1,0.3,0.6,0.6]$
$r_{\textsc{H-LLM}}=[0.4,0.0,0.3,0.0]$
Strictly Better: $L_{h}^{\text{opt}}=0.43$
$r_{\textsc{H-Only}}=[0.8,0.6,0.8,0.6]$
$r_{\textsc{H-LLM}}=[0.6,0.3,0.7,0.1]$
Strictly Better: $L_{h}^{\text{opt}}=0.09$
$r_{\textsc{H-Only}}=[0.3,0.9,0.3,0.9]$
$r_{\textsc{H-LLM}}=[0.0,0.1,0.0,0.3]$
Algorithm $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$
H-Only $0.30\pm 0.03$ $0.0$ $0.41\pm 0.04$ $0.0$ $0.68\pm 0.03$ $0.0$ $0.54\pm 0.05$ $0.0$
H-LLM $0.21\pm 0.03$ $0.1$ $0.16\pm 0.04$ $0.1$ $0.42\pm 0.06$ $0.1$ $0.09\pm 0.02$ $0.1$
Population $0.13\pm 0.03$ $0.05\pm 0.01$ $0.16\pm 0.04$ $0.05\pm 0.01$ $0.53\pm 0.05$ $0.05\pm 0.01$ $0.35\pm 0.07$ $0.05\pm 0.01$
Modiste-LinUCB $0.15\pm 0.05$ $0.04\pm 0.01$ $\mathbf{0.15\pm 0.04}$ $\mathbf{0.06\pm 0.01}$ $0.51\pm 0.06$ $0.06\pm 0.01$ $0.20\pm 0.10$ $0.06\pm 0.01$
Modiste-KNN $\mathbf{0.17\pm 0.03}$ $\mathbf{0.03\pm 0.01}$ $0.15\pm 0.04$ $0.07\pm 0.01$ $\mathbf{0.46\pm 0.05}$ $\mathbf{0.05\pm 0.02}$ $\mathbf{0.12\pm 0.03}$ $\mathbf{0.08\pm 0.01}$
Computational Validation
To validate Modiste computationally, we need access to simulated human behavior. Decision-makers may have different “expertise” (i.e., strengths and weaknesses) across input space $\mathcal{X}$ under each form of support. To construct realistic synthetic decision-makers, we set up a pilot study to understand these differences. This allows us to see how expertise varies across decision-makers and identify the personalized policy we can expect Modiste to learn.

### v3

@@ Abstract
Individual human decision-makers may benefit from different forms of support to improve decision outcomes, but when will each form of support yield better outcomes? In this work, we posit that personalizing access to decision support tools can be an effective mechanism for instantiating the appropriate use of AI assistance. Specifically, we propose the general problem of learning a decision support policy that, for a given input, chooses which form of support to provide to decision-makers for whom we initially have no prior information. We develop Modiste, an interactive tool to learn personalized decision support policies. Modiste leverages stochastic contextual bandit techniques to personalize a decision support policy for each decision-maker. In our computational experiments, we characterize the expertise profiles of decision-makers for whom personalized policies will outperform offline policies, including population-wide baselines. Our experiments include realistic forms of support (e.g., expert consensus and predictions from a large language model) on vision and language tasks. Our human subject experiments add nuance to and bolster our computational experiments, demonstrating the practical utility of personalized policies when real users benefit from accessing support across tasks.
1 Introduction
@@
Human decision-makers use various forms of support to inform their opinions before making a final decision (Keen 1980). Decision-makers with differing expertise may benefit from different forms of support on a given input (Yu et al. 2024). For example, one radiologist may provide a better diagnosis of a chest X-ray by leveraging model predictions (Kahn Jr 1994) while another may perform better after viewing suggestions from senior radiologists (Briggs et al. 2008): see Figure 1. In this paper, we study how to improve decision outcomes by personalizing which form of support we provide to a decision-maker on a case-by-case basis.
@@
Figure 1: Depending on the input, decision-makers need different forms of decision support to make correct decisions. Modiste personalizes access to the right form of support at the right time for the right decision-maker online. Here, Alice would not benefit from model access, while Bob would not benefit from a senior consult.
@@
Since artificial intelligence (AI) is increasingly used as a form of decision support (Lai et al. 2023), even moving towards systems that could act as “thought partners” (Collins et al. 2024b), responsible machine learning (ML) model deployment requires clarity on who should have access to model outputs and when model outputs can be safely exposed to decision-makers (Amodei et al. 2016). Regulation increasingly calls for the “effective and appropriate use” of AI (Biden 2023), requiring careful consideration of when models ought to be accessible to decision-makers. In this paper, we formalize learning a decision support policy that dictates for each individual decision-maker when additional support (e.g., LLM output) should be viewed and used for a given input.
@@
While prior work has assumed access to offline human decisions under support (Laidlaw and Russell 2021; Charusaie et al. 2022) or oracle queries of human behavior (De-Arteaga, Dubrawski, and Chouldechova 2018; Mozannar and Sontag 2020) to learn decision support policies, we argue that this data is unrealistic to obtain in practice across all available forms of support for a new decision-maker. Thus, for individuals for whom we have no prior information initially, we propose learning how to personalize support online. We develop Modiste, [^1] While a “modiste” usually refers to someone who tailors clothing and makes dresses/hats, we use the term to capture our tool’s ability to alter a policy to a decision-maker. an interactive tool that leverages off-the-shelf stochastic contextual bandit algorithms  (Li et al. 2010) to learn decision support policies by modeling human prediction error under varying forms of support.
@@
Our computational experiments explore the utility of personalization across multiple expertise profiles. Based on these experiments, we characterize decision-maker expertise into profiles where personalized policies outperform offline ones, such as population-wide majority vote. We demonstrate that if there is no benefit of personalization, Modiste recovers the same performance as the best form of support.
@@
To validate Modiste on real users ($N=80$), we conduct human subject experiments, where we explore forms of support that include expert consensus, outputs from an LLM, or predictions from a classification model. In contrast to prior work that only tests offline policies or evaluates in simulation, we demonstrate how Modiste can be used to learn personalized decision support policies online on both vision and language tasks. We emphasize our main contributions:
@@
1. Formalizing decision support policies. We propose a formulation for learning a personalized decision support policy that selects the form of support that maximizes a given decision-maker’s performance. We introduce Modiste, a tool to instantiate our formulation using existing methods from stochastic contextual bandits to model human prediction error under different forms of support. We open-source Modiste as a tool to encourage the adoption of personalized decision support policies.
@@
2. Evaluating personalized policies in realistic settings. We use Modiste to learn personalized policies for new decision-makers through both computational and human subject experiments on vision and language tasks. We characterize under which settings we would expect personalized policies to improve performance. Our human subject experiments, where real users interact with Modiste, nuance our findings from the computational experiments on synthetic decision-makers, demonstrating the appropriate use of decision support has benefits in practice.
2 Related Work
@@
Our study of decision support policies has implications for safely deploying ML models to interact with users. This topic is of heightened importance, particularly in light of recent calls for the “effective and appropriate use” of AI in US President Biden’s Executive Order (Biden 2023) and for clarity on when to “decide not to use [an] AI system” per the EU AI Act (EUA 2023). The disuse of AI can caution downstream misuse of models for assistive decision-making (Brundage et al. 2018). The refusal to use AI assistance can be strategic to empower decision-makers, thus preventing their overreliance on models and encouraging their agency on the task at hand (Gordon and Mugar 2020; Barabas 2022). Each decision-maker may require a different level of use to promote effective use of AI assistance in their decision-making (Kirk et al. 2024); for instance, experts and novices may prefer LLM access in different settings on theorem proving tasks (Collins et al. 2024a). Our experiments engage with such settings by learning when to provide LLM support for language-based tasks via a personalized policy.
@@
While various forms of decision support have been proposed, such as expert consensus (Scheife et al. 2015) and changes to machine interfaces (Roda 2011), more recent forms of support focus on algorithmic tools where decision-makers are aided by machine learning (ML) models (Phillips-Wren 2012; Gao et al. 2021; Bastani, Bastani, and Sinchaisri 2022). In some prior work, the human does not always make the final decision, such as those that learn to defer decisions from a model to a single decision-maker (Madras, Pitassi, and Zemel 2018; Mozannar and Sontag 2020) or others that jointly learn an allocation function between a model and a pool of decision makers (Keswani, Lease, and Kenthapadi 2021; Hemmer et al. 2022). In our setting, the human is always the decision-maker, which includes settings where humans make the final decisions with support from ML models (Green and Chen 2019; Lai et al. 2023), as well as those where humans make decisions when provided with additional information beyond a model prediction, e.g., explanations (Bansal et al. 2021b), uncertainty (Zhang, Liao, and Bellamy 2020), conformal sets (Babbar, Bhatt, and Weller 2022). While these studies always show a single form of support, recent works consider adapting when AI support is shown to users with a fixed policy. Ma et al. (2023) fit a decision tree to offline user’s decisions to decide when to show AI support to users, and Buçinca et al. (2024) use offline reinforcement learning to estimate if AI support would be helpful, especially under time constraints (Swaroop et al. 2024). Our work considers general forms of support, beyond when to show AI support, and formalizes learning in which contexts each form of support should be provided to an unseen decision-maker online. An extensive comparison to prior work is in the Appendix.
@@
Figure 2: We illustrate the process of learning a decision support policy $\pi_{t}$ online to improve a decision-maker $h$’s performance. Since assuming access to sufficient amounts of offline data is unreasonable in practice, our formulation learns a personalized policy online; each decision-maker’s learned policy may differ from that of another decision-maker if they have decisions ($\tilde{y}$) and thus different expertise.
3 Preliminaries
@@
4 Modiste: Learning Personalized Decision Support Policies
@@
To model the decision-making process of a human decision-maker without access to their previous decisions, we consider a stochastic contextual bandit set-up, where the forms of support are the arms, ${\mathcal{X}}$ is the context space, and the policy $\pi$ can be learned online. In Algorithm 1, we detail an algorithm for learning a policy $\pi^{*}$ that minimizes expected loss online. Our goal is to find an optimal decision support policy $\pi^{*}$ that minimizes $L_{h}(\pi)$. We can rewrite Eq. 1 as follows:
@@
where $r_{\textsc{A}_{i}}(x;h)=\mathbb{E}_{y|x}[\ell(y,h(x,\textsc{A}_{i}))]$ is the human prediction error for input $x$ and support $\textsc{A}_{i}$. Then, it can be shown that the optimal policy takes the form $\pi^{*}(x)=\operatorname*{arg\,min}_{\textsc{A}_{i}\in\mathcal{A}}r_{\textsc{A}_{i}}(x;h)$: see derivation in Appendix. For Modiste to run, we must maintain and update our estimate of human prediction error $r_{\textsc{A}_{i}}(x;h)$ (Step 1 in Algorithm 1) and of our policy $\pi$ (Step 2 in Algorithm 1).
@@
where $b_{\textsc{A}_{i},t}(x;h)$ corresponds to some exploration bonus. In the Appendix, we provide implementations of Modiste with LinUCB and with KNN.
@@
We provide an extendable interface for the study and deployment of decision-support policies. At each time step, Modiste sends each user’s predictions to a server running Algorithm 1, which identifies the next form of support for the next input. Modiste then updates the interface accordingly to reflect the selected form of support. Our tool can be flexiblely linked to crowdsourcing platforms like Prolific (Palan and Schitter 2018). We implement three common forms of support: (1) Human Alone, where the human makes the decision solely based on the input, (2) Model Prediction, which shows decision-makers a model’s prediction for the given input (Bastani, Bastani, and Sinchaisri 2022), and (3) Expert Consensus, which presents the user with a distribution over labels from multiple annotators (Scheife et al. 2015). In Figure 3, we provide an example screenshot of the interface. Participants are informed of their own correctness after each trial and the correctness of the form of support (e.g., model prediction) if support was provided, so that participants can learn whether support ought to be relied upon.
Figure 3: Example of the Modiste interface for MMLU-$2A$ where the human is provided responses from a LLM.
5 Experimental Set-up
Before we evaluate Modiste, we overview the set-up of subsequent experiments, both computational and human subject. All other details are in the Appendix.
@@ 5 Experimental Set-up
In this task, we consider three forms of support: Human Alone, Model, Consensus. Our goal is to construct a setup reflecting a realistic setting in which different forms of support result in different strengths and weaknesses for decision-makers. To instantiate this setting, we deliberately corrupt images of different classes to evoke performance differences – necessitating that a decision-maker appropriately calibrate when to rely on each form of support. We consider $5$ of the animal classes in CIFAR-10; of these, we never corrupt images of Birds, do not corrupt images of Deers and Cats for the Model, and do not corrupt images of Horses and Frogs for the Consensus.
@@
The two forms of support are Human Alone and LLM, where the human is provided responses generated from InstructGPT3.5, text-davinci-003(Ouyang et al. 2022) using the same few-shot prompting scheme for MMLU as Hendrycks et al. (2020). We conducted pilot studies to select a subset of topics where the accuracy of the LLM and average human accuracy vary. We choose the following topics: Computer Science, US Foreign Policy, High School Biology, and Elementary Mathematics. The goal of this task is to evaluate whether we can learn personalized support “in-the-wild,” where we naturally expect people to excel at different topics, akin to real-world settings where decision-makers may have varying expertise.
@@
  • Population-wide, where the decision-maker receives a form of support based on the majority vote from 10 learned policies (breaking ties at random). For this baseline, the form of support may vary across contexts but is not personalized to individual needs. This baseline is akin to recent offline policy learning (Ma et al. 2023; Buçinca et al. 2024).
@@
Modiste-KNN $0.01\pm 0.01$ $0.05\pm 0.03$ $\mathbf{0.05\pm 0.03}$
Table 1: We evaluate Modiste across three expertise profiles. We compute the average excess loss $L_{h}(\pi)-L_{h}^{opt}$ (lower is better), and standard deviation across individuals in each expertise profile for both CIFAR-$3A$ (Left) and MMLU-$2A$ (Right). $L_{h}(\pi)$ is computed by averaging across the last 10 steps of 100 total time steps. We bold the variant with the lowest excess loss that is statistically significant from the other variants. Note that this only occurs in the “varying” expertise setting.
6 Computational Evaluation
Decision-makers may have different “expertise” (i.e., strengths and weaknesses) across input space $\mathcal{X}$ under each form of support. We evaluate Modiste using simulated human behavior to capture diverse decision-maker expertise.